### v1

# LLMs Struggle to Perform Counterfactual Reasoning with Parametric Knowledge

Abstract
Large Language Models have been shown to contain extensive world knowledge in their parameters, enabling impressive performance on many knowledge intensive tasks. However, when deployed in novel settings, LLMs often encounter situations where they must integrate parametric knowledge with new or unfamiliar information. In this work, we explore whether LLMs can combine knowledge in-context with their parametric knowledge through the lens of counterfactual reasoning. Through synthetic and real experiments in multi-hop reasoning problems, we show that LLMs generally struggle with counterfactual reasoning, often resorting to exclusively using their parametric knowledge. Moreover, we show that simple post-hoc finetuning can struggle to instill counterfactual reasoning ability – often leading to degradation in stored parametric knowledge. Ultimately, our work reveals important limitations of current LLM’s abilities to re-purpose parametric knowledge in novel settings.
1 Introduction
Large Language Models (LLMs) internalize extensive world knowledge during pretraining, powering tasks like open-domain QA, fact retrieval, and knowledge-base completion (Petroni et al., 2019; Liu et al., 2019; Roberts et al., 2020). Benchmarks such as NaturalQuestions and HotpotQA focus on recall and multi-hop composition but do not require integrating novel premises at inference time (Yang et al., 2018; Kwiatkowski et al., 2019). For example:
“If Paris were located in Italy, in which country would the Eiffel Tower stand?”
Answering this demands two capabilities: Contextual Override—suppressing “Paris is in France”—and Selective Retrieval—retaining the Paris–Eiffel Tower link. Standard benchmarks do not probe this dual requirement.
Prior work on retrieval-augmented generation illustrates how external documents can both aid and confuse LLMs when facts disagree (Lewis et al., 2021), and studies of grokking transformers show multi-hop chaining without dynamic premise integration (Wang et al., 2024a). However, none systematically evaluate counterfactual multi-hop reasoning against a pretrained knowledge graph.
We ask: Can modern LLMs combine stored knowledge with in-context counterfactual premises to answer multi-hop queries correctly?
To address this, we introduce synthetic graph-based and real-world causal benchmarks that isolate reinforcing, additive, contradictory, and irrelevant contexts; empirically analyze GPT-4o to identify context-ignoring and context-overfitting failure modes under standard, chain-of-thought, and fine-tuned prompts; demonstrate that simple post-hoc fine-tuning yields only marginal gains on counterfactual tasks and can degrade performance on factual benchmarks; and discuss implications for interactive systems, retrieval-augmented pipelines, and safety-critical applications requiring accurate conditional reasoning.
Our results reveal a fundamental limitation: despite powerful fact memorization and retrieval, current LLMs lack mechanisms to dynamically modify their internal knowledge graph in response to new or conflicting information. Addressing this gap will require novel modeling and training paradigms.
2 Related Works
Multi-Hop QA
Multi-hop QA benchmarks measure a model’s ability to chain stored facts but do not test integration of new premises. Yang et al. (2018) introduce a dataset for contextual multi-hop QA, and Wang et al. (2024a) show that transformers can chain relations in a synthetic “grokking” regime. Real-world analyses compare reasoning over different relation types (Yang et al., 2024) and investigate how intermediate entities emerge in transformer layers (Biran et al., 2024). In contrast, we examine how LLMs combine parametric memory with in-context information.
Knowledge Conflicts
LLMs face knowledge conflicts when external context contradicts internal facts. Techniques like contextual decoding constraints (Yuan et al., 2024) and attention pruning (Li et al., 2025) enforce premises, while closed-book QA relies solely on retrieved parameters (Petroni et al., 2019; Roberts et al., 2020). Retrieval-augmented methods (REALM (Guu et al., 2020), RAG (Lewis et al., 2021), DPR (Karpukhin et al., 2020), FiD (Izacard & Grave, 2020)) merge external evidence, and finer-grained approaches like AdaCAD (Wang et al., 2024b) and CD2 (Jin et al., 2024) balance parametric and contextual sources. However, none target selective retention of stored knowledge under counterfactual premises.
Causal Reasoning and Counterfactuals in NLP
Counterfactual reasoning with LLMs remains challenging. Benchmarks such as CLadder (Jin et al., 2023), QRData (Liu et al., 2024), and CounterBench (Chen et al., 2025) reveal poor performance on formal causal and counterfactual tasks. Critics argue LLMs often “parrot” causal patterns from data rather than infer causally (Zečević et al., 2023), and Yamin et al. (2024) detail failures like over-reliance on parametric knowledge. While methods exist for generating faithful explanations (Gat et al., 2024) or counterfactual tokens (Chatzi et al., 2024), a systematic study of multi-hop counterfactual integration is still lacking.
3 Experiments on Real-World LLMs
3.1 Problem Formulation for a Causal Case
We inject contextual information at test time. For example, if the model’s parametric knowledge includes and prompt asserts , can it deduce ? Results shown in main paper are for GPT 4o (OpenAI et al., 2024), and LLama 3.1 8B (Grattafiori et al., 2024) results can be found in the Appendix. We categorize contextual edits into four partitions (all prompts and code are in the appendix/supplement):
1.
Scenario 1 (Reinforcing Prior Knowledge): Prompting the LLM with a relationship already present in its prior knowledge graph, thereby reinforcing an existing edge. Example: Given excessive rain causes flooding, query whether excessive rain causes infrastructure damage.
2.
Scenario 2 (Adding New Information): Prompting the LLM with scenario-specific information necessary to answer the query, but absent from its parametric knowledge graph, akin to adding an edge. Example: Informing the LLM that excessive rain causes Timmy to eat vegetables, and querying whether excessive rain improves Timmy’s health.
3.
Scenario 3 (Contradicting Prior Knowledge): Prompting the LLM with information that strongly contradicts its existing parametric knowledge, equivalent to replacing an edge in its prior knowledge graph. Example: Informing the LLM that excessive rain causes desert expansion and querying whether excessive rain promotes cactus growth.
4.
Scenario 4 (Irrelevant Information): Prompting the LLM with unrelated information, akin to providing an edge from a disconnected knowledge graph. Example: Informing the LLM that fatty food consumption causes heart attacks, and querying whether excessive rain leads to flooding.
3.2 Prompting Methods
We compare three strategies: Standard—direct causal query; CoT—chain-of-thought prompting; FT—fine-tuning on counterfactual examples with CoT explanations (160 examples, hyperparameters in Appendix). Models are queried over either 1 or two counterfactual hops for simplicity.
3.3 Results
We ask a binary cause/non-cause question (random baseline 50). Figure 1 reports GPT-4o accuracy across scenarios.
3.3.1 Success when Context Does Not Oppose Prior
We see that in Scenario 1 where we reinforce prior knowledge, Figure 1(a) shows that standard prompting achieves around 85 accuracy, with CoT accuracy rises above 95, and FT further closes errors to near 100 with low variance across trials. These results together demonstrate that when contextual information reinforces existing knowledge, modern LLMs like GPT-4o can reliably utilize their parametric knowledge without being misled by the prompt. Such robustness provides a useful baseline against which to compare the more challenging counterfactual scenarios in Scenarios 2 and 3.
3.3.2 Failure when Context Adds New Information or Contradicts Prior
In the adding-information scenario 2 plotted in Figure 1(b), CoT yields 80 accuracy (30 above random) but shows variability across relation types; FT improves this to 85 by reinforcing task-specific patterns. Under conflicting premises plotted in Figure 1(c), CoT performance collapses to near the 50 baseline, with responses oscillating between stored and contextual facts. Fine-tuning partially mitigates this, lifting accuracy to 75, yet significant errors persist, highlighting the difficulty of overriding strong parametric priors. As such, it becomes clear that the greatest hurdle we encounter is information that conflicts with our prior. In a sense, scenario 2 , the second worst performance, where we add new information can be viewed as a weaker version of prior conflicting information. For example, if we look back at the previous example where we inform the LLM that excessive rain causes Timmy to eat vegetables, the LLM likely has a weak prior that rain does not cause kids to eat vegetables in general.
3.4 Mixed Results for Irrelevant Information
In the irrelevant-information scenario 4 plotted in Figure 1(b), standard prompts receive around 65 accuracy, while CoT recovers to 75 and FT boosts performance to 90. It should be noted that LLama 3.1 8B results (appendix) show finetuning reduces performance. As such, we see mixed signals in this regime, possibly owing to the large difference in parameter sizes in these models.
4 Conceptual Experiments in Toy Setting
In the previous section, we demonstrated that state-of-the-art LLMs can struggle to perform counterfactual reasoning tasks and that simple fine-tuning can be insufficient to overcome this. In this section, we perform experiments in a more controlled setting to better understand the origins of this limitation. This setting can be viewed as a generalization of the causal inference setting studied in the previous section – where previously we considered only a single relation (of causality).
4.1 Setup
We adapt Wang et al. (2024a)’s synthetic knowledge graph (entities , relations ). Facts are Atomic or Multi-Hop paths. Atomic Facts represent only a single edge in the knowledge graph (i.e. ent1 r1 ent2). On the other hand, Multi-Hop Facts represent a composition of two atomic facts (i.e. ent1 r1 ent2 r2 ent3). Pretraining follows Wang et al. (2024a) (including transformer layer setup), ensuring greater than 99% validation accuracy on multi-hop composition. We perform the pretraining stage using the hyper-parameters reported in Wang et al. (2024a). We extend this to counterfactual multi-hop QA: given a novel atomic counterfactual premise, answer a multi-hop query assuming it’s true, testing integration with parametric knowledge.
Counterfactual Finetuning
For simplicity, we concentrate our work on two-hop settings. We then have three possible types of counterfactual multi-hop queries. Hop 1-Relevant are those in which the counterfactual premise supplied is relevant to the first hop of the multi-hop reasoning problem. Hop 2-Relevant are those in which the counterfactual premise is relevant to the second hop of the multi-hop reasoning problem. Both of these are classified as Related queries. Finally, Unrelated queries are those in which the counterfactual premise should not effect the outcome of the final query. We ensure that the counterfactual fine-tuning dataset is balanced across these three catagories to mitigate shortcut learning.
4.2 Findings
Counterfactual Finetuning Induces Shortcuts
Figure 2(a) shows that, although finetuning quickly drives accuracy on relevant counterfactual queries to high levels, performance on irrelevant queries remains poor—even when training on a balanced mix. Our results point to the learning of a shortcut solution whereby the model becomes induced to always override its contextual knowledge even when the counterfactual premise is irrelevant. We additionally examine the sensitivity of this behavior to finetuning hyper-parameters in Figure 2(b), finding that it arises across a range of learning rates.
Performance Degradations Not a Result of Format Change
One potential explanation for the difficulty in introducing counterfactual reasoning could be the mismatch in format between the original presentation of knowledge and the counterfactual style prompts. To isolate the impact of prompt format, we finetuned on an “Unrelated Only” dataset (Figure 2(c)), where counterfactual premises never affect the correct answer—so the model only needs to adapt to a new prompt style while relying on its existing parametric knowledge. Achieving near perfect accuracy shows that format changes (or generic forgetting) don’t drive the drop in performance; instead, the core difficulty lies in teaching the LLM to conditionally override its parametric knowledge without corrupting it.
Effect of Incorporating Counterfactual Data in Pretraining
We explored adding counterfactual data to pretraining (augmenting 20% of KG edges, Figure 2d). Incorporating counterfactual data indistinguishably from other pretraining data (Augmented in plot) or marking it with a special token (Marked-Augmented in plot) in pretraining induce good accuracies across the different groups of counterfactual points (i.e unrelated and related). Marked-Augmented performances converges faster than the Augmented case. However, while we find that interleaving counterfactual data into pretraining induces counterfactual reasoning, regular fact learning is slowed as can be seen in Figure 3 where we plot factual Marked-Augmented performances . This suggests a tradeoff between learning counterfactual reasoning and facts at pretraining-time. It is possible that counterfactual tasks introduce interfering gradients which contribute to the suppression of parametric knowledge.
5 Discussion
Our work highlights LLM challenges in counterfactual reasoning requiring dynamic integration of parametric and contextual knowledge. Experiments show LLMs often default to parametric knowledge or fail to balance contextual override with selective retrieval. Simple finetuning is limited, inducing shortcuts or degrading knowledge. Pretraining with counterfactual data, while improving such reasoning, can also harm factual task performance. These findings point to a core limitation: current LLMs lack robust mechanisms for on-the-fly, conditional use of their parametric knowledge.
6 Limitations
While our study sheds light on the challenges LLMs face in integrating parametric knowledge with novel counterfactual premises, it is subject to several limitations. Our synthetic benchmarks abstract away many complexities of real-world reasoning. In the toy setting, counterfactual premises are expressed as single-edge edits to a static knowledge graph and queries are limited to two-hop chains. Many real-world scenarios often involve multi-predicate interactions, ambiguous or probabilistic relationships, and noisy or conflicting evidence from multiple sources. What is significant is that we show how models struggle even under these simpler circumstances.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix
All code is in supplementary materials
Appendix A Causal Case
A.1 Full Prompt Library
Below are every prompt template employed in our experiments. Placeholders (‘<…>‘) were dynamically filled in the code with the appropriate event names or lists, as described below.
Placeholder Definitions:
‘<Cause>‘: The event assumed to be the cause in the causal-query prompt (e.g., “Rain”).
‘<Effect>‘: The event assumed to be the effect in the causal-query prompt (e.g., “Flooding”).
‘<Query>‘: The event whose causal relationship is being queried (e.g., “Crop Growth”).
‘<Target>‘: The event we are checking if ‘<Query>‘ causes (e.g., “Food Supply Increase”).
‘<PrevEvent>‘: The most recently generated event in the chain when expanding synthetic event sequences.
‘<ExistingList>‘: A serialized list of all previously generated events in that chain (as a Python list string), used to prevent duplicates.
A.2 A. General Causal-Query Prompts
1. Standard (Direct) Prompt
Assume that <Cause> causes <Effect>. Given this statement, does <Query> logically cause <Target>? Only output your answer as <answer>Yes/No</answer> Feel free to extrapolate beyond the given context.
2. Chain-of-Thought (CoT) Prompt
Assume that <Cause> causes <Effect>. Please think step by step to decide whether <Query> logically causes <Target>. Then output your reasoning and your answer: <answer>Yes/No</answer>
> **Fine-Tuned** model runs used these exact templates but pointed at the fine-tuned checkpoint.
A.3 Event Chain Generation Prompts
1. Causal Chain Expansion
generate an event that is caused by the event: <PrevEvent> (if this event is empty pick a random choice, do not simply ever output the text "if this event is empty pick a random choice"), for example Cancer -> Death or Obesity -> Bad Heart Health. Use an arrow between the two events such that <PrevEvent> is the first item in the chain. Make sure the event you generate is not already in the list: <ExistingList>. Make sure the output only includes the two events with an arrow between them.
2. Anticausal Chain Expansion
generate an event that is anticausal to <PrevEvent> (meaning the effect is actually the opposite of what it should be), for example Cancer -> Longer Life or Obesity -> Weight Loss. Use an arrow between the two events. Make sure the event you generate is not already in the list: <ExistingList>. Output only the two events with an arrow between them.
3. “Irrelevant” Transition Event
generate a random event that is a result of <PrevEvent> (meaning this is a specific strange scenario), for example Rain -> Increased Chocolate Eating or Obesity -> Warm Weather. This event should not typically be a result of <PrevEvent>. Use an arrow between the two events. Start with <PrevEvent> and separate the events with an ->.
4. Post-Transition Chain Expansion
generate an event that is caused by the event: <PrevEvent> (if this event is empty pick a random choice), for example Cancer -> Death or Obesity -> Bad Heart Health. Use an arrow between the two events such that <PrevEvent> is the first item in the chain. Make sure the event you generate is not already in the list: <ExistingList>. Output only the two events with an -> between them.
Appendix B Llama Results
In these results in Figure 4 , we see similar trends to the GPT-4o results except that we see Fine Tuning producing decreased accuracy for irrelevant counterfactuals. This mirrors what we see in the toy example.
Appendix C GPT Finetuning HyperParameters
Trained tokens: 38,754 Epochs: 3 Batch size: 1 LR multiplier: 2
Appendix D LLama 3.1 8B Finetuning HyperParameters
Epochs: 5, LoRA Rank:8,Learning Rate: 0.0001, Max Context Length:8192
Appendix E LLM Usage
Our paper focuses on examining the reasoning abilities of Language Models. We do not use a language model to assist with writing.
Appendix F Compute
In our toy experiment, we used 4 gpus per experiment. The GPU used was NVIDIA A6000 and we had a total of 72 GPU hours.

### v2

@@
-LLMs Struggle to Perform Counterfactual Reasoning with Parametric Knowledge
+Can LLMs Reconcile Knowledge Conflicts in Counterfactual Reasoning?
 Abstract
@@
 1 Introduction
-Large Language Models (LLMs) internalize extensive world knowledge during pretraining, powering tasks like open-domain QA, fact retrieval, and knowledge-base completion (Petroni et al., 2019; Liu et al., 2019; Roberts et al., 2020).
-Benchmarks such as NaturalQuestions and HotpotQA focus on recall and multi-hop composition but do not require integrating novel premises at inference time (Yang et al., 2018; Kwiatkowski et al., 2019).
-For example:
+Large Language Models (LLMs) internalize vast amounts of world knowledge during pretraining, enabling impressive performance on a wide range of knowledge‐intensive tasks such as open‐domain question answering, fact retrieval, and knowledge‐base completion (Petroni et al., 2019; Liu et al., 2019; Roberts et al., 2020).
+Benchmarks like NaturalQuestions and HotpotQA have driven progress on recall‐based and multi‐hop reasoning, but they primarily evaluate a model’s ability to regurgitate stored facts or compose chains of parametric knowledge without new external inputs (Yang et al., 2018; Kwiatkowski et al., 2019).
+In contrast, many real‐world scenarios require LLMs to integrate their pretrained knowledge with novel or hypothetical information provided at inference time.
+For example, consider a counterfactual query:
 “If Paris were located in Italy, in which country would the Eiffel Tower stand?”
-Answering this demands two capabilities: Contextual Override—suppressing “Paris is in France”—and Selective Retrieval—retaining the Paris–Eiffel Tower link.
-Standard benchmarks do not probe this dual requirement.
-Prior work on retrieval-augmented generation illustrates how external documents can both aid and confuse LLMs when facts disagree (Lewis et al., 2021), and studies of grokking transformers show multi-hop chaining without dynamic premise integration (Wang et al., 2024a).
-However, none systematically evaluate counterfactual multi-hop reasoning against a pretrained knowledge graph.
-We ask: Can modern LLMs combine stored knowledge with in-context counterfactual premises to answer multi-hop queries correctly?
-To address this, we introduce synthetic graph-based and real-world causal benchmarks that isolate reinforcing, additive, contradictory, and irrelevant contexts; empirically analyze GPT-4o to identify context-ignoring and context-overfitting failure modes under standard, chain-of-thought, and fine-tuned prompts; demonstrate that simple post-hoc fine-tuning yields only marginal gains on counterfactual tasks and can degrade performance on factual benchmarks; and discuss implications for interactive systems, retrieval-augmented pipelines, and safety-critical applications requiring accurate conditional reasoning.
-Our results reveal a fundamental limitation: despite powerful fact memorization and retrieval, current LLMs lack mechanisms to dynamically modify their internal knowledge graph in response to new or conflicting information.
-Addressing this gap will require novel modeling and training paradigms.
+Answering this correctly demands two distinct capabilities: Contextual Override and Selective Retrieval.
+In order to override parametric knowledge using context, the model must temporarily suppress its default fact that “Paris is in France” and accept the hypothetical premise.
+Furthermore, the model must still retrieve and leverage the association between “Eiffel Tower” and “Paris” stored in its weights, even though the location of Paris has been altered.
+Standard QA and multi‐hop benchmarks do not explicitly test this dual requirement.
+A growing body of work has examined knowledge conflicts and context‐based overrides in LLMs.
+Studies on retrieval‐augmented generation highlight how external documents can both help and confuse a model when facts disagree (Lewis et al., 2021).
+Recent analyses of multi‐hop reasoning in synthetic settings (“grokking” transformers) show that models can learn to chain stored relations but typically without the capacity to absorb new premises on the fly (Wang et al., 2024a).
+Yet none of these works systematically probe counterfactual multi‐hop reasoning, where the premise may conflict with or extend the pretrained knowledge graph.
+In this paper, we fill that gap by asking: Can modern LLMs selectively combine parametric knowledge with in‐context counterfactual premises to answer multi‐hop questions correctly?
+Our key contributions are:
+Counterfactual QA benchmarks.
+We introduce both synthetic graph‐based tasks (extending “Grokked” reasoning benchmarks) and real‐world causal reasoning scenarios to isolate cases of (i) reinforcing, (ii) adding, (iii) contradicting, and (iv) irrelevant context relative to the pretrained knowledge graph.
+Empirical analysis.
+Through experiments on GPT‐4o and other state‐of‐the‐art models, we identify two primary failure modes: (a) context‐ignoring (model defaults to stored facts) and (b) context‐overfitting (model blindly follows the prompt).
+We quantify performance across standard, chain‐of‐thought (CoT), and fine‐tuned prompting strategies.
+Fine‐tuning pitfalls.
+We demonstrate that simple post‐hoc fine‐tuning on counterfactual examples often yields only marginal gains on the target tasks and can degrade performance on standard factual benchmarks by inducing unintended heuristics.
+Practical implications.
+We discuss the impact of our findings for interactive systems, retrieval‐augmented pipelines, and safety‐critical applications where accurate conditional reasoning under novel premises is essential.
+Taken together, our results reveal a fundamental limitation of current LLMs: despite their remarkable capacity to memorize and retrieve facts, they lack robust mechanisms for on‐the‐fly modification or extension of their internal knowledge graph in response to conflicting or new information.
+Addressing this gap will require new modeling and training paradigms that can dynamically integrate stored and contextual knowledge without compromising either.
 2 Related Works
 Multi-Hop QA
-Multi-hop QA benchmarks measure a model’s ability to chain stored facts but do not test integration of new premises.
+Multi-hop question answering has been examined in a variety of prior works as a measure of knowledge manipulation ability.
+(Yang et al., 2018) introduces a benchmark for measuring multi-hop question-answering abilities for context-based question-answering.
+Allen-Zhu & Li (2024) examines knowledge manipulation in controlled settings, finding that LLMs can struggle to solve one-step knowledge manipulation tasks without COT or extensive fine-tuning.
+Wang et al.
+(2024a), on the other hand, shows that transformers can implicitly solve multi-hop queries in the grokking regime on a synthetic knowledge-graph task.
+Abramov et al.
+(2025) extends these findings to more real settings by augmenting Wikipedia data.
 Yang et al.
-(2018) introduce a dataset for contextual multi-hop QA, and Wang et al.
-(2024a) show that transformers can chain relations in a synthetic “grokking” regime.
-Real-world analyses compare reasoning over different relation types (Yang et al., 2024) and investigate how intermediate entities emerge in transformer layers (Biran et al., 2024).
-In contrast, we examine how LLMs combine parametric memory with in-context information.
+(2024b) study the reliability of latent multi-hop reasoning on real-world knowledge, finding that certain types of relations are more conducive to multi-hop reasoning than others.
+Biran et al.
+(2024) studies the mechanistic implementation of multi-hop question-answering in LLMs, finding that it tends to arise when intermediate entities can be resolved in earlier layers of the model.
+.
+Unlike prior works, which primarily focus on multi-hop reasoning involving only parametric knowledge, we study cases in which an LLM may need to combine parametric and knowledge from its context.
 Knowledge Conflicts
-LLMs face knowledge conflicts when external context contradicts internal facts.
-Techniques like contextual decoding constraints (Yuan et al., 2024) and attention pruning (Li et al., 2025) enforce premises, while closed-book QA relies solely on retrieved parameters (Petroni et al., 2019; Roberts et al., 2020).
-Retrieval-augmented methods (REALM (Guu et al., 2020), RAG (Lewis et al., 2021), DPR (Karpukhin et al., 2020), FiD (Izacard & Grave, 2020)) merge external evidence, and finer-grained approaches like AdaCAD (Wang et al., 2024b) and CD2 (Jin et al., 2024) balance parametric and contextual sources.
-However, none target selective retention of stored knowledge under counterfactual premises.
+Large Language Models (LLMs) face challenges with knowledge conflicts, where external context clashes with internal parametric knowledge.
+Responses often prioritize context, parametric stores, or blend them for factual accuracy.
+Some methods enforce contextual premises (Yuan et al., 2024) or use attention pruning for context-exclusive outputs (Li et al., 2025).
+While useful for overriding facts, these approaches may be ill-suited for counterfactual reasoning, which requires selective retention and integration of parametric knowledge, not its wholesale dismissal.
+Conversely, closed-book QA relies solely on stored parametric facts (Petroni et al., 2019; Roberts et al., 2020) but cannot adapt to novel information or reason under hypothetical conditions deviating from this knowledge.
+Hybrid techniques, especially retrieval-augmented generation (RAG), attempt to merge parametric memory with retrieved information.
+Methods like REALM (Guu et al., 2020), RAG (Lewis et al., 2021), Dense Passage Retrieval (DPR) (Karpukhin et al., 2020), and FiD (Izacard & Grave, 2020) condition generation on external evidence, while advanced strategies such as AdaCAD (Wang et al., 2024b) and CD2 (Jin et al., 2024) offer finer-grained balancing of these knowledge sources.
+Goyal et al.
+(2025) studies the dynamics of finetuning models for context reliance, finding that instruction tuning can often worsen context reliance.
+Although these methods enhance the incorporation of external factual information, they are generally not designed for the distinct challenges of counterfactual reasoning.
+Specifically, they do not enable LLMs to accept a hypothetical premise contradicting parametric facts, then selectively retrieve relevant parametric knowledge, and subsequently perform complex multi-hop reasoning based on this newly integrated understanding.
 Causal Reasoning and Counterfactuals in NLP
-Counterfactual reasoning with LLMs remains challenging.
-Benchmarks such as CLadder (Jin et al., 2023), QRData (Liu et al., 2024), and CounterBench (Chen et al., 2025) reveal poor performance on formal causal and counterfactual tasks.
-Critics argue LLMs often “parrot” causal patterns from data rather than infer causally (Zečević et al., 2023), and Yamin et al.
-(2024) detail failures like over-reliance on parametric knowledge.
-While methods exist for generating faithful explanations (Gat et al., 2024) or counterfactual tokens (Chatzi et al., 2024), a systematic study of multi-hop counterfactual integration is still lacking.
+Counterfactual reasoning is increasingly vital in Natural Language Processing (NLP), especially with Large Language Models (LLMs), whose causal reasoning capabilities are an active research area (Liu et al., 2024b).
+While some propose LLMs as a new frontier for textual causal discovery (Kiciman et al., 2024), critical views suggest they might be "causal parrots" merely echoing training data (Zečević et al., 2023).
+Benchmarking efforts consistently reveal LLM limitations: CLadder assesses formal causal reasoning (Jin et al., 2023), QRData highlights struggles with data-based causal tasks (Liu et al., 2024a), and CounterBench shows poor performance on formal counterfactual inference (Chen et al., 2025).
+Studies like Yamin et al.
+(2024) further detail these issues, identifying LLM failure modes such as over-reliance on parametric knowledge or narrative shortcuts.
+Despite these struggles, research actively explores using LLMs for specific NLP counterfactual tasks, like generating faithful explanations (Gat et al., 2024) or alternative textual outputs via counterfactual token generation (Chatzi et al., 2024).
+These endeavors underscore a critical gap: a deep understanding of how and why LLMs falter when integrating their vast parametric knowledge with novel, in-context counterfactual premises, particularly in multi-hop reasoning.
 3 Experiments on Real-World LLMs
 3.1 Problem Formulation for a Causal Case
-We inject contextual information at test time.
-For example, if the model’s parametric knowledge includes and prompt asserts , can it deduce ?
-Results shown in main paper are for GPT 4o (OpenAI et al., 2024), and LLama 3.1 8B (Grattafiori et al., 2024) results can be found in the Appendix.
-We categorize contextual edits into four partitions (all prompts and code are in the appendix/supplement):
+We start off by testing the LLM’s capacity to reason about a very simple type of relationship between events – basic causality.
+In this example, we task the LLM with the binary question of determining if one event causes another event.
+Let’s assume we have in the LLM’s parametric (pre-training) knowledge that where are events that occur in the real world such as "rain falling" or "plants growing." We use the notation to indicate direct causation.
+Simply asking the LLM if is a cause of would constitute a one-hop query, and asking the LLM if causes constitutes a two-hop query.
+We note that state of the art LLMs like GPT 4o generally succeed at these kinds of tasks.
+(Yang et al., 2024a) (All code and prompts for entire paper are in Appendix/Supplementary Materials).
+Now suppose that we further introduce contextual information at test-time.
+For example, what if the LLM is told that instead of , we have that such that the LLM knows in its parametric knowledge that .
+Can the LLM deduce that is now a cause of ?
+Now we have to deal with a more complex counterfactual multi-hop scenario.
+In the main body of the paper, we include results from GPT-4o, Open AI’s SOTA reasoning model GPT-5 (Thinking), a fine-tuned version of GPT-4o (GPT-5 is not fine-tuneable) and include results from LLama 3.1 in the Appendix.
+We categorize the contextual information introduced to the LLM into 4 partitions, as follows:
 1.
@@
 We compare three strategies: Standard—direct causal query; CoT—chain-of-thought prompting; FT—fine-tuning on counterfactual examples with CoT explanations (160 examples, hyperparameters in Appendix).
-Models are queried over either 1 or two counterfactual hops for simplicity.
+Models are queried over either 1 or 2 counterfactual hops for simplicity.
 3.3 Results
-We ask a binary cause/non-cause question (random baseline 50). Figure 1 reports GPT-4o accuracy across scenarios.
+We ask a binary cause/non-cause question (random baseline 50). Figure 2 reports GPT-4o and GPT-5 accuracy across the 4 scenarios.
 3.3.1 Success when Context Does Not Oppose Prior
-We see that in Scenario 1 where we reinforce prior knowledge, Figure 1(a) shows that standard prompting achieves around 85 accuracy, with CoT accuracy rises above 95, and FT further closes errors to near 100 with low variance across trials.
-These results together demonstrate that when contextual information reinforces existing knowledge, modern LLMs like GPT-4o can reliably utilize their parametric knowledge without being misled by the prompt.
+We see that in Scenario 1 where we reinforce prior knowledge, Figure 2(a) shows that standard prompting on GPT-4o with and without CoT, GPT-5(Thinking) and GPT-4o FT, the models perform very well with results ranging between 90 accuracy and perfect.
+These results together demonstrate that when contextual information reinforces existing knowledge, modern LLMs like GPT-4o and GPT-5 can reliably utilize their parametric knowledge without being misled by the prompt.
 Such robustness provides a useful baseline against which to compare the more challenging counterfactual scenarios in Scenarios 2 and 3.
 3.3.2 Failure when Context Adds New Information or Contradicts Prior
-In the adding-information scenario 2 plotted in Figure 1(b), CoT yields 80 accuracy (30 above random) but shows variability across relation types; FT improves this to 85 by reinforcing task-specific patterns.
-Under conflicting premises plotted in Figure 1(c), CoT performance collapses to near the 50 baseline, with responses oscillating between stored and contextual facts.
-Fine-tuning partially mitigates this, lifting accuracy to 75, yet significant errors persist, highlighting the difficulty of overriding strong parametric priors.
+In the adding-information scenario 2 plotted in Figure 2(b), the non fine-tuned models show performance rating between 60 and 75 accuracy while FT improves this to 90 by reinforcing task-specific patterns.
+Under conflicting premises plotted in Figure 2(c), performance collapses to near the 50 baseline with responses oscillating between stored and contextual facts with fine-tuning only marginally improving accuracy.
+As significant errors persist even with finetuning, this highlights the difficulty of overriding strong parametric priors.
 As such, it becomes clear that the greatest hurdle we encounter is information that conflicts with our prior.
@@
 For example, if we look back at the previous example where we inform the LLM that excessive rain causes Timmy to eat vegetables, the LLM likely has a weak prior that rain does not cause kids to eat vegetables in general.
-3.4 Mixed Results for Irrelevant Information
-In the irrelevant-information scenario 4 plotted in Figure 1(b), standard prompts receive around 65 accuracy, while CoT recovers to 75 and FT boosts performance to 90.
+3.3.3 Mixed Results for Irrelevant Information
+In the irrelevant-information scenario 4 plotted in Figure 2(d), we see strong performance across the board with finetuning on GPT 4o increasing the accuracy over standard and CoT GPT 4o prompting, while the reasoning model GPT-5 achieves near perfect accuracy.
 It should be noted that LLama 3.1 8B results (appendix) show finetuning reduces performance.
 As such, we see mixed signals in this regime, possibly owing to the large difference in parameter sizes in these models.
+3.4 Alignment
+Modern production LLMs are typically subjected to additional factuality and safety alignment stages—e.g., instruction tuning and reinforcement learning from human (or AI) feedback—which explicitly reward consistency with memorized facts and discourage speculative deviations (Ouyang et al., 2022; Stiennon et al., 2020; Bai et al., 2022a; b; Glaese et al., 2022).
+This post-training pressure biases models toward relying on their pretrained parametric knowledge even when prompts present plausible counterfactual premises, making contextual overrides harder to enact reliably.
+This alignment likely plays a part in the results we see.
+To disentangle this alignment from our results, we now train our own model in the coming section.
 4 Conceptual Experiments in Toy Setting
 In the previous section, we demonstrated that state-of-the-art LLMs can struggle to perform counterfactual reasoning tasks and that simple fine-tuning can be insufficient to overcome this.
-In this section, we perform experiments in a more controlled setting to better understand the origins of this limitation.
-This setting can be viewed as a generalization of the causal inference setting studied in the previous section – where previously we considered only a single relation (of causality).
+In this section, we perform experiments in a synthetic setup to better understand the mechanisms behind this difficulty.
+Our experiments take place on small transformers that we train from scratch, allowing us to study counterfactual reasoning capabilities independently of potential confounders arising from the factuality or alignment post-training of production LLMs.
 4.1 Setup
-We adapt Wang et al.
-(2024a)’s synthetic knowledge graph (entities , relations ).
-Facts are Atomic or Multi-Hop paths.
-Atomic Facts represent only a single edge in the knowledge graph (i.e.
-ent1 r1 ent2).
-On the other hand, Multi-Hop Facts represent a composition of two atomic facts (i.e.
-ent1 r1 ent2 r2 ent3).
-Pretraining follows Wang et al.
-(2024a) (including transformer layer setup), ensuring greater than 99% validation accuracy on multi-hop composition.
-We perform the pretraining stage using the hyper-parameters reported in Wang et al.
+We perform controlled experiments in a synthetic knowledge graph setting, as used in prior studies of knowledge-based reasoning Wang et al.
 (2024a).
-We extend this to counterfactual multi-hop QA: given a novel atomic counterfactual premise, answer a multi-hop query assuming it’s true, testing integration with parametric knowledge.
-Counterfactual Finetuning
-For simplicity, we concentrate our work on two-hop settings.
-We then have three possible types of counterfactual multi-hop queries.
-Hop 1-Relevant are those in which the counterfactual premise supplied is relevant to the first hop of the multi-hop reasoning problem.
-Hop 2-Relevant are those in which the counterfactual premise is relevant to the second hop of the multi-hop reasoning problem.
-Both of these are classified as Related queries.
-Finally, Unrelated queries are those in which the counterfactual premise should not effect the outcome of the final query.
-We ensure that the counterfactual fine-tuning dataset is balanced across these three catagories to mitigate shortcut learning.
+To summarize, we randomly generate a directed graph , where each vertex represents an entity from a set of entities where entities are linked by relations .
+For simplicity, each entity and relation is identified by a unique token in the synthetic language that we study.
+The graph structure induces two forms of knowledge: atomic facts and inferred facts.
+Atomic facts describe a single edge in the knowledge graph and can be expressed as a triple of . These represent basic relationships between entities that must be memorized by the model.
+Inferred Facts are those that can be deduced by a two-hop composition of the atomic facts.
+Inferred facts can be denoted as , where there exists such that .
+In the inferred fact setting, we refer to the first entity, , as the head entity and the final entity, , as the tail.
+The intermediate entity, , is referred to as the “bridge" entity.
+We refer to the atomic fact as the first hop and as the second hop.
+Intuitively, all atomic facts must be explicitly seen in order to specify the knowledge graph.
+On the other hand, a model with compositional capabilities can potentially generalize to predicting unseen inferred facts.
+Knowledge Graph Pretraining
+We follow the training setup proposed in Wang et al.
+(2024a) to train a model that is capable of deducing unseen inferred facts by composition.
+Concretely, the model is presented with all atomic facts and a subset of the inferred facts.
+The remainder of inferred facts are reserved for the test set.
+We perform the pretraining for many epochs, until the model is capable of deducing the unseen inferred facts.
+Full knowledge pretraining hyperparameters are provided in the Appendix and largely mirror the settings reported in Wang et al.
+(2024a).
+Counterfactual Reasoning Task
+Our primary interest in this work is the ability of LLMs to perform counterfactual reasoning over their internalized knowledge.
+To study this, we propose the following counterfactual reasoning task which requires the model to selectively contextually override its parametric knowledge with respect to certain entities, while selectively retrieving parametric knowledge relative to other entities.
+We consider the setting in which the model must incorporate a new relation from the context and combine this contextual knowledge with existing parametric knowledge.
+Concretely, we study a setting where a modified relation, which we term as the counterfactual premise is input in the context, along with an inferred fact query.
+The LLM must then answer the inferred fact as if the contextually provided relation was in the knowledge graph.
+As inferred facts require composing two relations, successfully performing the counterfactual reasoning task requires the LLM to simultaneously override its parameteric knowledge (to incorporate the counterfactual premise) while also making use of its parametric knowledge for the remaining relation.
+We additionally consider cases where the counterfactual premise is irrelevant to the inferred fact query (i.e.
+does not involve either the head or the bridge entity).
+Finetuning Stage
+In our synthetic setup, the language model is only exposed to knowledge graph-consistent relations during pretraining.
+Mirroring our experiments on frontier model, we finetune these knowledge graph-pretrained models from the previous stage on the counterfactual reasoning task which relies partially on the pretrained parametric knowledge.
+As finetuning often occurs on a much narrower distribution than model pretraining, we restrict the head entities used in finetuning to come from a 20% subset of the total entities in the knowledge graph.
+We also balance the training and test sets equally between the following classes of examples:
+1.
+Hop 1 Relevant Counterfactual: Here, the counterfactual premise modifies the first hop of the inferred fact.
+As an example, consider the query “If Paris was in Spain, what language would be spoken in Paris?".
+Here, it is discernible directly from the prompt that the counterfactual premise should is relevant to the query.
+2.
+Hop 2 Relevant Counterfactual: Here the counterfactual premise modifies the link between the bridge entity and the final answer.
+As an example consider the query “if Paris were located in Italy, in which country would the Eiffel Tower stand?".
+This represents a more challenging instance of counterfactual reasoning as the model must perform the first hop before determining that the counterfactual query is relevant to the query.
+3.
+Irrelevant Counterfactual: Finally, there may be cases in which the counterfactual premise is entirely irrelevant to the multi-hop query.
+As an example, consider the multi-hop query “If New York were located in Canada, in which country would the Eiffel Tower stand?" Here, the presence of the counterfactual premise should logically not impact the final answer.
+Legend
+Incorporating Counterfactual Reasoning during Pretraining
+We also examine a setting where counterfactual queries are mixed in during the pretraining stage.
+Concretely, we mixed in the training counterfactual prompt originally used for finetuning during the first stage of training (when atomic and memorized facts are being learned).
+We considered two settings: Augmented in which the counterfactual examples are added in without modification and Marked-Augmented in which the counterfactual examples are tagged with a special token before and added in the pretraining stage.
 4.2 Findings
 Counterfactual Finetuning Induces Shortcuts
-Figure 2(a) shows that, although finetuning quickly drives accuracy on relevant counterfactual queries to high levels, performance on irrelevant queries remains poor—even when training on a balanced mix.
-Our results point to the learning of a shortcut solution whereby the model becomes induced to always override its contextual knowledge even when the counterfactual premise is irrelevant.
-We additionally examine the sensitivity of this behavior to finetuning hyper-parameters in Figure 2(b), finding that it arises across a range of learning rates.
+In Figure 4(a),4(b) and 4(c) we show the performance of the model across the three splits throughout the finetuning process.
+Broadly, our results highlight that finetuning on the counterfactual reasoning task is highly susceptible to learning various shortcuts and, as a result, fails to perform uniformly well across the three splits of the evaluation dataset.
+Concretely, in the first stage of training, the model quickly learns to simply repeat the entity shown in the counterfactual premise.
+This leads to high accuracy in the Hop 2-Relevant split (Figure 4(b)) (where the final answer is always the entity present in the counterfactual premise in the query).
+As training continues, on the other hand, the model’s performance worsens on the Hop 2-Relevant split , and the model transitions from performing badly to performing well on the Hop 1-Relevant split (Figure 4(a)).
+Notably, the model never achieves strong performance on the Irrelevant counterfactual split, despite this split being equally represented in the finetuning training set (Figure 4(c)).
+Our findings suggest a key difficulty on this task lies in learning the selective override mechanism: the model is capable of combining contextual and parametric knowledge, but is unable to learn to distinguish when the counterfactual premise is relevant.
 Performance Degradations Not a Result of Format Change
-One potential explanation for the difficulty in introducing counterfactual reasoning could be the mismatch in format between the original presentation of knowledge and the counterfactual style prompts.
-To isolate the impact of prompt format, we finetuned on an “Unrelated Only” dataset (Figure 2(c)), where counterfactual premises never affect the correct answer—so the model only needs to adapt to a new prompt style while relying on its existing parametric knowledge.
-Achieving near perfect accuracy shows that format changes (or generic forgetting) don’t drive the drop in performance; instead, the core difficulty lies in teaching the LLM to conditionally override its parametric knowledge without corrupting it.
+One potential explanation for the observed challenge of incorporating counterfactual reasoning ability during finetuning could arise from a more generic phenomena of catastrophic forgetting by which the model may forget its pretraining knowledge as it is further trained.
+To control for performance degradations induced by mere fine-tuning, we constructed an additional downstream task which does not require any contextual override of pre-trained knowledge.
+Concretely, we induce a format change where, rather than directly output the response to the inferred fact query, the model must generate a chain-of-thought trace of the intermediate relations and entities from the head to the tail entities.
+As shown in Figure 4(d), fine-tuning is able to quickly adapt the model to this task, as evidenced by the 100% test accuracy.
+This suggests that the observed failure of fine-tuning arises not from generic catastrophic forgetting, but depends on the counterfactual reasoning task.
+We hypothesize that the training dynamics involved in contextual override can exacerbate catastrophic forgetting.
 Effect of Incorporating Counterfactual Data in Pretraining
-We explored adding counterfactual data to pretraining (augmenting 20% of KG edges, Figure 2d).
+We explored adding counterfactual data to pretraining (augmenting 20% of KG edges, Figure 5).
 Incorporating counterfactual data indistinguishably from other pretraining data (Augmented in plot) or marking it with a special token (Marked-Augmented in plot) in pretraining induce good accuracies across the different groups of counterfactual points (i.e unrelated and related).
 Marked-Augmented performances converges faster than the Augmented case.
-However, while we find that interleaving counterfactual data into pretraining induces counterfactual reasoning, regular fact learning is slowed as can be seen in Figure 3 where we plot factual Marked-Augmented performances .
-This suggests a tradeoff between learning counterfactual reasoning and facts at pretraining-time.
 It is possible that counterfactual tasks introduce interfering gradients which contribute to the suppression of parametric knowledge.
+To summarize, our findings in simulation highlight that it is challenging for LLMs to selectively override their parametric knowledge during pretraining.
+In the synthetic setting, counterfactual finetuning quickly pushes the model towards over-relying on context.
+Notably, this contrasts with the behavior we observe in production models, where performance primarily suffers on conflicting knowledge scenarios.
+We attribute this difference to the fact that production models are likely to undergo significant factuality training which could bias their behavior towards consistency with parametric knowledge.
+Taken together our findings demonstrate that performing counterfactual reasoning tasks requires a delicate balance between preserving and overriding parametric knowledge, which is difficult to instill during finetuning.
 5 Discussion
@@
 Pretraining with counterfactual data, while improving such reasoning, can also harm factual task performance.
-These findings point to a core limitation: current LLMs lack robust mechanisms for on-the-fly, conditional use of their parametric knowledge.
-6 Limitations
+These findings point to a core limitation: current LLMs lack robust mechanisms for on-the-fly, conditional use of their internal knowledge.
+In practice, we observe two recurring failure modes—context-ignoring and context-overfitting—that persist across prompting strategies and fine-tuning.
 While our study sheds light on the challenges LLMs face in integrating parametric knowledge with novel counterfactual premises, it is subject to several limitations.
-Our synthetic benchmarks abstract away many complexities of real-world reasoning.
-In the toy setting, counterfactual premises are expressed as single-edge edits to a static knowledge graph and queries are limited to two-hop chains.
-Many real-world scenarios often involve multi-predicate interactions, ambiguous or probabilistic relationships, and noisy or conflicting evidence from multiple sources.
+In the toy setting, counterfactual premises are expressed as single‐edge edits to a static knowledge graph and queries are limited to two‐hop chains.
+Many real-world scenarios often involve multi‐predicate interactions, ambiguous or probabilistic relationships, and noisy or conflicting evidence from multiple sources.
 What is significant is that we show how models struggle even under these simpler circumstances.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Future work could focus on extending our analysis to deeper and more noisy multi-hop relations.
 Appendix
@@
 Placeholder Definitions:
-‘<Cause>‘: The event assumed to be the cause in the causal-query prompt (e.g., “Rain”).
-‘<Effect>‘: The event assumed to be the effect in the causal-query prompt (e.g., “Flooding”).
+‘<Cause>‘: The event assumed to be the cause in the causal‐query prompt (e.g., “Rain”).
+‘<Effect>‘: The event assumed to be the effect in the causal‐query prompt (e.g., “Flooding”).
 ‘<Query>‘: The event whose causal relationship is being queried (e.g., “Crop Growth”).
@@
 ‘<ExistingList>‘: A serialized list of all previously generated events in that chain (as a Python list string), used to prevent duplicates.
-A.2 A. General Causal-Query Prompts
+A.2 A. General Causal‐Query Prompts
 1. Standard (Direct) Prompt
 Assume that <Cause> causes <Effect>. Given this statement, does <Query> logically cause <Target>? Only output your answer as <answer>Yes/No</answer> Feel free to extrapolate beyond the given context.
-2. Chain-of-Thought (CoT) Prompt
+2. Chain‐of‐Thought (CoT) Prompt
 Assume that <Cause> causes <Effect>. Please think step by step to decide whether <Query> logically causes <Target>. Then output your reasoning and your answer: <answer>Yes/No</answer>
-> **Fine-Tuned** model runs used these exact templates but pointed at the fine-tuned checkpoint.
+> **Fine‐Tuned** model runs used these exact templates but pointed at the fine‐tuned checkpoint.
 A.3 Event Chain Generation Prompts
@@
 Start with <PrevEvent> and separate the events with an ->.
-4. Post-Transition Chain Expansion
+4. Post‐Transition Chain Expansion
 generate an event that is caused by the event: <PrevEvent> (if this event is empty pick a random choice), for example Cancer -> Death or Obesity -> Bad Heart Health.
@@
 Our paper focuses on examining the reasoning abilities of Language Models. We do not use a language model to assist with writing.
-Appendix F Compute
+Appendix F Toy Experiment Hyperparameters
+| Hyperparameter | Values |
+| Learning Rate | |
+| Weight Decay | |
+| Batch Size | 512 |
+Appendix G Compute
 In our toy experiment, we used 4 gpus per experiment. The GPU used was NVIDIA A6000 and we had a total of 72 GPU hours.